Exhibit 22-13

Quantity of Output

(Q)

Total Fixed Cost (TFC)

Average Fixed Cost (AFC)

Total Variable Cost
(TVC)

Average Variable Cost
(AVC)

Total Cost
(TC)

Average Total Cost
(ATC)

Marginal Cost
(MC)

0

$200

$0

$200

1

$200

(A)

30

(H)

230

(M)

(S)

2

$200

(B)

50

(I)

250

(N)

(T)

3

$200

(C)

(F)

$26.67

(K)

(P)

(U)

4

$200

(D)

130

(J)

330

(Q)

(V)

5

$200

(E)

(G)

$40

(L)

(R)

(W)

Refer to Exhibit 22-13.  What dollar amounts go in blanks (I) and (J), respectively?

Group of answer choices

$100; $50

$25; $32.50

$500; $40

$50; $32

There is not enough information to answer this question.
Answer #1

Ans) the correct option is b) $25; $32.50

Average variable cost = Variable cost / output

I = 50/2 = 25

J = 130/4 = 32.50